UT 27 November 2018

Split night with Eng (1st half) / AZ (2nd half)

Night Center 00:07 (Local Time) - switch will be at 00:02 (local time) to accomodate reconfig from LUCI to MODS

OSA: S. Allanson

ISp: J. Morris

ISA: B. Rothberg (Tucson)

Software Support. P. Kubank (Tucson)

AO Support: A. Vaz (2nd half - Tucson), G. Taylor (Tucson)

Additional Support: J. Hill, O. Kuhn (Tucson)

Overview:

Engineering night - obatined ESM vs SL data for highly variable seeing conditions over 3 targets across three filters. One set (M52) we obtained observations with bright (R=10) and dim (R=15.3) AO ref stars. Able to close the loop on a double AO Ref star for Palomar 2. Conditions were as close to photometric as possible. Opportunites for testing GCS updates. Petr and John H. (with assistance from Doug M.) seemed to make a good amount of progress dealing with GCS issues. We had no occurrences of co-pointing errors with SL+ESM offsets to sky. Steve A. operated ESM for 3 targets, across various AO Ref star brightness with no significant problems. Handover occurred at precisely 00:02 (reconfiguration time was split between LBTO Eng and AZ). One LUCI-2 MOS error occurred early on (downtime ~5 min).

AZ Science - successful observations of SN target in NGC 1068 for PI Sand. Obtained MODS-1 Dual grating science data with high S/N in continuum. Seeing variable. Reconfigured for LUCI-2 AO observations for PI Ertel. Obtained a significant amount of data for target. Ran through the script 3x plus a few extra exposures.  ADI and collimation scripts worked great.  PI was pleased.


UT 12:37 - LUCIs software up and running, LUCI-1 mask put away, both LUCIs initialized, attempted to wake both MODS, MODS-1 happy, but MODS-2 Blue was cross-fibered with MODS-2 Red (slight scare on my part) - J Riedl will fix before we need to use it.

UT 12:38 - starting at pointing star until time to point and collimate

UT 12:51 - collimating

UT 12:56 - AO checkout is done

UT 12:58 - LUCI-2 MOS error

UT 01:04 - LUCI-2 MOS recovered - (both SGs were on, but then SG2 turned off on its own - ran init MOS after conifrming Mask in FPU, then commanded mask to be put away)

L1 0127-0128 - 2.51 sec dark test exposures

L2 0001-0002 - 2.51 sec dark test exposures

UT 01:10 - sending L1/SL + L2/ESM script to M51

UT 01:12 - LUCI-2 FW error - bringing up eng panel - inititalized - and made manual move to Clear

UT 01:14 - Steve doing manual acquire for AO/LUCI-2

UT 01:18 - RunAO happy, skipping past FW1-L2 error and beginning exposures

L2 FM4 is off a lot, will manually fix after first exposure

L2 -0004 - verification on FM4 move - looks OK

UT 01:22 - offset to sky - worked

UT 01:25 - DIMM 1.1" SX WFS 1.1"

UT 01:36 - network drop out

UT 01:40 - H-band now DIMM at 1.36", SX WFS 1.0"

UT 01:58 - J-band now DIMM 1.57" SX WFS 1.2"

UT 02:07 - had to restart RTD2 - image only half-appeared

L1 - 0129 - 0161

L2 - 0003 (skip 0004) - 0036

Resend similar preset to M52 but for fainter AO Ref star (15.2 R)

UT 02:17 - sending bino preset SL/LUCI-1 & ESM/LUCI-2 M52

UT 02:18 - guidsetar missed on SX

UT 02:19 - resending preset - DIMM 1.4"

UT 02:24 - active optics ran away on right (DX) - Steve working on it

UT 02:25 - restarting GCS right

UT 02:27 - resending preset

UT 02:31 - DX is old version of GCS - SX is new version - resending preset again

UT 02:33 - Petr testing GCS on right - putting new one in again to check runaway, resending preset

UT 02:35 - cleared active optics on DX, resending preset, testing new version of DX to see if we get runaway again - order of magnitude error it seems (3000 vs 300 etc for zernikes)

UT 02:37 - putting old version of GCS back in for DX

UT 02:38 - resending yet again

UT 02:40 - faint AO Ref star- center stars not getting it, Steve manually adjusting it

UT 02:43 - optimizing ESM

UT 02:47 - starting data collection

UT 03:07 - H-band Fainter AORef star DIMM 1.33"

L1 0162-0175 K complete, H 3/11 exposures done

L2 0037-0050 K complete, H 3/11 exposures done

UT 03:10 - Updating GCS patch will stop and restart DX

UT 03:11 - resending preset

UT 03:14 - active optics running away on both sides

UT 03:15 - reverting back GCS, resending preset

UT 09:18 - DIMM 1.58" - see overdensity of noise - Steve centering up there (R=15.3)

UT 03:19 - will restart with H-band

UT 03:21 - optimizing

L1 0176-0186 H-band only

L2 0051-0061 H-band only

L1 186 - blocking zernikes being sent - J. Hill testing issues with GCS WFS

UT 03:44 - stopping GCS - changing to new version to test

UT 03:45 - resending preset - will pick up with J-band exposures

preset failed

UT 03:46 - preset resent - failed

UT 03:49 - resending preset - DIMM 1.8"

UT 03:55 - need to restart CCD39 process

UT 03:55 - resending preset - conked out again - Steve calling Greg

UT 04:02 - too many CCD39 viewere caused problems

UT 04:04 - GCS left was stuck on pause, although ESM was nice and happy, resending preset - will skip to J-band part of script

UT 04:13 - loops closed and happy - skipping to J-band

UT 04:22 - DIMM back down to 1"

L1 0187-0197 J-band

L2 0062-0062 J-band

UT 04:32 script complete, John Hill restarting GCS on left

UT 04:34 - doing a pointing check first (blind filter in LUCIs)

Going to do Palomar 2 to get more data and SL data

UT 04:39 -sending Pal2 preset with brighter AORef star (that is actually a double star)

UT 04:40 - interrupted system call on AOS - meaning it will fail, it did, resent preset

UT 04:48 - taking data - we'll see if double star is ok or if we get weird PSFs - DIMM = 0.91" SX WFS 0.9" -looks good

UT 04:52 - SX (SL) images look elongated - only in upper part of pupil

UT 05:04 - DIMM 1.2" SX WFS 1.19"

UT 05:52 - DIMM and WFS has gotten sub-1" - seeing 0.6" on ESM side

UT 05:56 - completed script

L1 0198-0236

L2 0063-0111

UT 05:58 - sent FS6 binocular script (Standard Star)

UT 06:01 - missed it on DX pointing

-pointing check

UT 06:04 - sending FS6 (Feige 22) again - airmass 1.14

UT 06:06 - mountain drop out communications

L1 0237-0248

L2 0112-0123

UT 06:24 - moving to HCG31 - ESM test

UT 06:32 - started exposures

UT 06:45 - DX loops didn't close on return to field from off - this is the first ESM hiccup of the night

UT 06:47 - Steve recovered - but we got a RIP

UT 06:49 - quick pointing check needed

L1 0249-0257

L2 0124-0132 - not a complete K - but enough for a nice pic/analysis

UT 06:51 - resending preset

UT 06:58 - contiuing script in an on position


UT 07:02 - switching to MODS-1, put LUCIs in safe state

UT 07:07 - mountain connection dropout

UT 07:09 - mountain connection dropout again

UT 07:18 - moving to pointing star

UT 07:20 - collimation check

UT 07:26 - DIMM 0.93"

UT 07:29 - initial collimation is struggling - Steve working on it

UT 07:32 - sending PI Sands preset (AZ) to SN2018ivc in NGC 1068

UT 07:39 - correcting alignment by -0.25" DETX, nudging back in by +0.13 DETX

mods1r.20181127.0001-0005 - aquisition and alignment

UT 07:42 - starting spectroscopy

UT 07:45 - exposures are 120sec x 5 exposures - will check first exposures to see if one can go longer without saturating

first exposures show at most 600 counts in SN continuum (host galaxy h-alpha reaches maybe 1000 counts) - will bump to 1200seconds

mods1r.20181127.0006-0007 - 120 seconds

mods1b.20181127.0001-0002 - 120 seconds

mods1r.20181127.0008 - 1200 seconds

mods1b.20181127.0003 - 1200 seconds

DIMM jumped up to 1.42" - had been ~1"

UT 08:15 - communications dropout again from mountain

UT 08:16 - finished SAND program, will move to SpecStd

UT 08:17 - sending G191B2b acquisition

UT 08:19 - weird error - acquisition script looked like it passed a BLUE PARTNER BLUE PI info but that was not in the script. CTRLC'd it and tried again, preset went through

mods1r.20181127.0009 - g191b2b alignment

mods1r.20181127.0010 - thru slit g191b2b

used modsalign -r to tweak offset in thru slit

UT 08:26 - starting spectroscopic exposures

mods1r.20181127.0011-0013 - G191b2b

mods1b.20181127.0004-0006 - G191b2b

UT 08:40 - completed MODS-1 observations


UT 08:40 - reconfiguring for LUCI-Bino

Running init all again on LUCIs

UT 08:45 - authorizing LUCI-2 DX only

UT 08:47 - putting N30 fieldstop into LUCI-2 FPU

UT 08:53 - pointing check

UT 08:54 - sending collimation preset

UT 08:57 - sending target script WD0912+536

UT 09:00 - DIMM 1.17"

L2 0133 - 2.51sec blind exposure

L2 0134 - 5sec H-band image N375 (had to skip to row 2 to force it to setup for this exposure - not sure why) - matched finder chart

L2 0135- 10sec H-band N30FS in FPU w/N375 camera

star is at 1054,1069 - offseting -3.5" -5.3' DETXY

L2 0136/0137 - test exposure to see if centered - 10sec x 2NDIT

Absorbing the offset

L2 0138 - field stop off - aligning

L2 0139 - test of FM4 alignment

UT 09:15 - Starting Science exposures

-keeping eye on FieldStop motion

UT 09:41 - DIMM now 1.3"

UT 09:45 - DIMM now 1.70"

UT 10:02 - DIMM now 1.27"

UT 10:06 - DIMM now 1.09"

UT 10:19 - Mountain drop out again

L2 0230 - jump during exposure - don't use

UT 10:21 - DIMM now 1.08"

UT 10:41 - DIMM now 0.98"

- FM4/Fieldstop alignment relatively stable

UT 11:06 - DIMM now 0.8"

UT 11:09 - script round 1 completed

UT 11:09 - restarting script from beginning - DIMM 1.09"

UT 11:49 - DIMM now 1.01"

UT 12:35 - plan is to go until we can't AO anymore (and before sun comes up)

UT 01:03 - finished 2nd round, starting a third round of script

UT 01:24 - starting to get too bright

UT 01:25 - will finish after 3rd block in round three of script

L2 0140-0299 - first round of script

L2 0300-0459 - second round of script

L2 0460-0489 -third round of script

UT 01:26 - putting LUCI-2 in safe config - closing dome for calibrations

Will take LUCI-1/LUCI-2 JHK N375 Flats and MODS-1 cals

UT 12:38 - starting arc lamps MODS-1 mods1b.20181127.0007-0030 ; mods1r.20181127.0014-0032

L1 0258-0262 lights off J

L1 0263-0267 lights on J

L1 0268-0272 lights off H

L1 0273-0277 lights on H

L1 0278-0282 lights off K

L1 0283-0287 lights on K

L2 0490-0494 lights off J

L2 0495-0499 lights on J

L2 0500-0504 lights off H

L2 0505-0509 lights on H

L2 0510-0514 lights off K

L2 0515-0519 lights on K

UT 01:52 - Done with LUCI-Bino cals - putting LUCI in safe config

UT 02:41 - put MODS-1 to sleep

ALL DONE!

-- BarryRothberg - 27 Nov 2018
Topic revision: r1 - 27 Nov 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback